OpenStudy (anonymous):

3 cube root 375?????

OpenStudy (anonymous):

\[\sqrt[3]{375}\] ???

OpenStudy (anonymous):

theres a 3 in front of that .

OpenStudy (anonymous):

\[3\sqrt[3]{375}\] ?

OpenStudy (anonymous):

So, you are looking for "x" such that x^3 = 375

OpenStudy (anonymous):

@gaara438125 yes

OpenStudy (anonymous):

Since cubic roots are very seldom taught anymore, I am thinking that you are just expected to use a calculator. There should be a button on your calculator that looks like:\[\sqrt[x]{y}\] And in this case x=3 and y=375. Type in 375 first, hit the button and then 3 and then "=".

OpenStudy (raden):